OpenStudy (anonymous):

A freshwater tank holds 400 gallons and evaporates 2 gallons per week. A salt water tank holds 451 gallons and evaporates 5 gallons per week. In how many weeks will the two tanks hold the same amount? I think its 17 weeks

OpenStudy (anonymous):

Let x be the number of weeks 400 - 2x = 451 -5x x = 17 17 weeks

OpenStudy (lilg132):

451 - 5x = 400 - 2x 451 - 400 = 5x - 2x 51 = 3x 17 = x